# Minimax Hypothesis Testing for the Bradley-Terry-Luce Model

### v1

@@ Abstract
[…]
Index Terms:
Bradley-Terry-Luce model, hypothesis testing, minimax risk, spectral methods, principal ratios of Markov matrices.
@@ I Introduction
In recent years, the availability of pairwise comparison data and its subsequent analysis has significantly increased across diverse domains. Pairwise comparison data consists of information gathered in the form of comparisons made among a given set of items or agents. Many real-world applications, including sports tournaments, consumer preference surveys, and political voting, generate data in the form of pairwise comparisons. Such datasets serve a range of purposes, such as ranking items [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12], analyzing team performance over time [13], studying market or sports competitiveness [14, 15], and even fine-tuning large language models using reinforcement learning from human feedback [16, 17].
A popular modeling assumption while performing such learning and inference tasks with pairwise comparison data is to assume that the data conforms to an underlying Bradley-Terry-Luce (BTL) model [2, 3, 4, 5, 6] as a generative model for the data. Given $n$ items $\{1,\dots,n\}$, the BTL model assigns a latent “skill score” $\alpha_{i}>0$ to each item $i$, representing its relative merit or utility compared to other items, and posits that the likelihood of $i$ being preferred over an item $j$ in a pairwise comparison is given by
@@
The BTL model is known to be a natural consequence of the assumption of independence of irrelevant alternatives (IIA), which is widely used in economics and social choice theory [3]. However, the IIA assumption underlying the BTL model has been questioned, and it has been shown that the BTL model may not always accurately capture real-world datasets [18, 19, 20]. For instance, in the case of sports tournaments, the BTL model is incapable of capturing the home-advantage effect, which refers to the possible advantage that a home-team may experience when playing against a visiting team. The home-advantage effect has been observed in several sports, such as soccer [21] and cricket [22], and ignoring this effect can lead to biased estimates of the skill scores of teams. Additionally, real-world datasets may exhibit non-transitive behaviors that violate the BTL model. To address these limitations, modifications of the BTL model that incorporate the home-advantage effect [23, Chapter 10], Thurstonian models [24], and other generalizations of the BTL model [7] have been explored in the literature. Furthermore, to capture more complex behaviors within pairwise comparison datasets, various other models have been introduced, such as models based on Borda scores [25, 26] and non-parametric stochastically transitive models [27, 28, 29].
@@
Nevertheless, primarily because of its simplicity and interpretability, the BTL framework remains one of the most widely used models. A large fraction of the associated results in the literature focuses on estimation of the skill score parameters of the BTL model. Some popular approaches include maximum likelihood estimation [6, 7], rank centrality (or Markov-chain-based) methods [30, 10], least-squares methods [12], and non-parametric methods [27, 13] (also see [8, 9] for Bayesian inference for BTL models). Once the parameters are estimated, they are then used for inference tasks, such as ranking items and learning skill distributions [14, 15]. As alluded to above, an inherent assumption for such statistical analysis to be meaningful in real-world scenarios is that the BTL model accurately represents the given pairwise comparison data. If the underlying data-generating process violates the assumptions made by the BTL model, then the model’s predictions and guarantees cannot be trusted. Hence, it is important to develop a systematic method to test the validity of the BTL assumption on real-world datasets. Such a method can help identify scenarios where a BTL model can provide a useful approximation to the underlying preferences or skills, thereby providing essential guidance for downstream applications.
[…]
@@ I-A Related Literature
This work lies at the confluence of two fields of study: preference learning and hypothesis testing. On the preference learning front, the problem of analyzing ranked preference data has become increasingly important with the rise in the availability of data on consumer preferences and web surveys. In a complementary vein, the analysis of preference data, such as pairwise comparisons, has a rich history starting from the seminal works [2, 4, 3, 5, 24, 6]. Among the various models proposed for ranking and analyzing pairwise comparison data, the BTL model [2] has emerged as one of the most widely used and well-studied approaches. Initially introduced in [6] for estimating participants’ skill levels in chess tournaments, the BTL model is a special case of the more general Plackett-Luce model [3, 4], originally developed in social choice theory, mathematical psychology, and statistics contexts. The BTL model finds applications in diverse domains, such as sports [31, 32], psychology [33], ranking of journals [34], fine-tuning large language models [16], etc. We refer readers to [35, 36] for a comprehensive overview of different models of rankings.
@@
In the literature, many studies have focused on estimating parameters of the BTL model and characterizing the corresponding error bounds, cf. [37, 38, 7, 30, 10, 39, 28, 11]. For example, [10] introduces and analyzes spectral estimation methods, [11] presents non-asymptotic bounds for relative $\ell^{\infty}$-estimation errors of normalized vectors of skill scores for both spectral and maximum likelihood estimators, and [28] provides graph-dependent $\ell^{2}$-estimation error bounds for the maximum likelihood estimator. Another emerging line of work includes uncertainty quantification for the estimated parameters, cf. [40, 41, 42, 43]. For example, [40] proves the asymptotic normality of the maximum likelihood estimator as well as the spectral estimator for BTL skill parameters. In a different vein, [14, 15] develop a technique for estimating the underlying skill distribution of agents participating in a tournament, introducing a Bayesian flavor to the BTL estimation problem.
@@
On the other hand, hypothesis testing also has a rich history in statistics, ranging from Pearson’s $\chi^{2}$-test [44] to non-parametric tests [45]. The classical literature develops some tests for the IIA assumption, e.g., [46, 47], but these tests are based on asymptotic $\chi^{2}$-approximations. Moreover, testing for BTL models has been explored far less than the estimation question in recent times. And to the best of our knowledge, no study has developed rigorous hypothesis tests with minimax analysis to determine the validity of the BTL assumption in the literature. The minimax perspective of hypothesis testing, which we specialize in our setting, was initially proposed by [48]. Among related works to ours, [49] analyzed two-sample testing on pairwise comparison data, and [50] derived lower bounds for testing the IIA assumption given preference data. The formulations in [49, 50] are different to our formulation. For example, we quantify the separation distance of a pairwise comparison model from the class of BTL models in terms of the Frobenius norm rather than the variant of total variation (TV) distance used in [50]. For the special case of pairwise comparisons under a complete induced observation graph, the lower bounds in [50] agree with ours in terms of the high-level scaling law of the critical threshold with respect to $n$ and $k$. However, their study does not focus on providing statistical tests and corresponding upper bounds as we do in this work. Finally, in the alternative setting of online preference learning, [51] studied the testing problem for various forms of stochastic transitivities and established upper and lower bounds on the sample complexity for testing.
@@
In this work, we also briefly explore the stability of rankings made under a BTL assumption when the underlying model does not conform to this assumption. Indeed, since the BTL assumption may not perfectly hold in various scenarios, it is crucial to examine the stability of rankings obtained under the BTL assumption. In the literature, [25] provided empirical evidence that the BTL assumption is not very robust to changes in the pairwise comparison matrix, and [52] proposed the Sync-Rank algorithm which treated the ranking problem as a group synchronization problem, thereby eliminating the dependence of the algorithm on the BTL assumption. Furthermore, recent work by [53] demonstrated the ineffectiveness of standard spectral methods for BTL parameter estimation in the presence of even small fractions of Byzantine voters, and proposed a more robust Byzantine spectral ranking algorithm. In contrast to these works, we introduce and utilize a new measure of separation distance that quantifies deviation from the class of BTL models. As we will see, this measure is instrumental in our analysis to quantify the stability of BTL assumptions in the context of rankings with respect to classical Borda count rankings.
@@ I-B Main Contributions
[…]
@@
    We also prove several auxiliary results. For example, we utilize the notion of separation distance mentioned above to analyze the stability of BTL assumptions in the context of rankings. More specifically, we investigate the deviation from the BTL condition that is sufficient for the ranking produced under the BTL assumption to differ from the classical Borda count ranking [54]. Our results in Proposition 9 show that a deviation of $O(1/\sqrt{n})$ from the BTL condition is sufficient to produce inconsistent BTL and Borda rankings. As another example, during our analysis, we also obtain $\ell^{2}$-estimation error bounds for (virtual) BTL parameters in Lemma 6 when the data is actually generated by a general pairwise comparison model as opposed to a BTL model. These bounds also highlight the robustness of the spectral ranking method under model mismatch.

### v2

@@ Abstract
The Bradley-Terry-Luce (BTL) model is one of the most widely used models for ranking a collection of items or agents based on pairwise comparisons among them. Specifically, given $n$ agents, the BTL model endows each agent $i$ with a latent skill score $\alpha_{i}>0$ and posits that the probability that agent $i$ is preferred over agent $j$ in a comparison is $\alpha_{i}/(\alpha_{i}+\alpha_{j})$. In this work, our objective is to formulate a hypothesis test that determines whether a given pairwise comparison dataset, with $k$ comparisons per pair of agents, originates from an underlying BTL model. We formalize this testing problem in the minimax sense and define the critical threshold of the problem. We then establish upper bounds on the critical threshold for general observation graphs and highlight their dependence on two fundamental structural properties: principal ratio and edge expansion. Additionally, we derive lower bounds on the critical threshold for the special case of complete induced graphs, thereby demonstrating that the critical threshold scales as ${\Theta}((nk)^{-1/2})$ in a minimax sense in this case. In particular, our test statistic for the upper bounds is based on a new approximation we derive for the separation distance between general pairwise comparison models and the class of BTL models. To further assess the performance of our statistical test, we prove upper bounds on conditional probabilities of error. Additionally, we derive several other auxiliary results over the course of our analysis, such as bounds on principal ratios of graphs, $\ell^{2}$-bounds on BTL parameter estimation under model mismatch, stability of rankings under the BTL model with small model mismatch, etc. Finally, we conduct several experiments on synthetic and real-world datasets to validate some of our theoretical results. Moreover, we also propose an approach based on permutation testing to determine the threshold of our test in a data-driven manner in these experiments.
@@ I Introduction
In recent years, the availability of pairwise comparison data and its subsequent analysis has significantly increased across diverse domains. Pairwise comparison data consists of information gathered in the form of comparisons made among a given set of items or agents. Many real-world applications, including sports tournaments, consumer preference surveys, and political voting, generate data in the form of pairwise comparisons. Such datasets serve a range of purposes, such as ranking items [BradleyTerry1952, Luce1959, Plackett1975, McFadden1973, Zermelo1929, Hunter2004, GuiverSnelson2009, CaronDoucet2012, NegahbanOhShah2017, Chenetal2019, HendrickxOlshevskySaligrama2020], analyzing team performance over time [BongLiShrotriyaRinaldo2020], studying market or sports competitiveness [JadbabaieMakurShah2020, JadbabaieMakurShah2020journal], and even fine-tuning large language models using reinforcement learning from human feedback [InstructGPT, DPO].
A popular modeling assumption while performing such learning and inference tasks with pairwise comparison data is to assume that the data conforms to an underlying Bradley-Terry-Luce (BTL) model [BradleyTerry1952, Luce1959, Plackett1975, McFadden1973, Zermelo1929] as a generative model for the data. Given $n$ items $\{1,\dots,n\}$, the BTL model assigns a latent “skill score” $\alpha_{i}>0$ to each item $i$, representing its relative merit or utility compared to other items, and posits that the likelihood of $i$ being preferred over an item $j$ in a pairwise comparison is given by
@@
The BTL model is known to be a natural consequence of the assumption of independence of irrelevant alternatives (IIA), which is widely used in economics and social choice theory [Luce1959]. However, the IIA assumption underlying the BTL model has been questioned, and it has been shown that the BTL model may not always accurately capture real-world datasets [DavidsonMarschak1959, McLaughlinLuce1965, Tversky1972]. For instance, in the case of sports tournaments, the BTL model is incapable of capturing the home advantage effect, which refers to the possible advantage that a home team may experience when playing against a visiting team. The home advantage effect has been observed in several sports, such as soccer [Clarke1995] and cricket [Morley2005], and ignoring this effect can lead to biased estimates of the skill scores of teams. Additionally, real-world datasets may exhibit non-transitive behaviors that violate the BTL model. To address these limitations, modifications of the BTL model that incorporate the home advantage effect [Agresti2002, Chapter 10], Thurstonian models [Thurstone1927], and other generalizations of the BTL model [Hunter2004] have been explored in the literature. Furthermore, to capture more complex behaviors within pairwise comparison datasets, various other models have been introduced, such as models based on Borda scores [ShahWainwright2018, HeckelShahRamchandranWainwright2019] and non-parametric stochastically transitive models [Chatterjee2015, Shahetal2016, Mania2018].
@@
Nevertheless, primarily because of its simplicity and interpretability, the BTL framework remains one of the most widely used models. A large fraction of the associated results in the literature focuses on estimation of the skill score parameters of the BTL model. Some popular approaches include maximum likelihood estimation [Zermelo1929, Hunter2004, YCOM], rank centrality (or Markov-chain-based) methods [NegahbanOhShah2012, NegahbanOhShah2017], least-squares methods [HendrickxOlshevskySaligrama2020], and non-parametric methods [Chatterjee2015, BongLiShrotriyaRinaldo2020] (also see [GuiverSnelson2009, CaronDoucet2012] for Bayesian inference for BTL models). Once the parameters are estimated, they are then used for inference tasks, such as ranking items and learning skill distributions [JadbabaieMakurShah2020, JadbabaieMakurShah2020journal]. As alluded to above, an inherent assumption for such statistical analysis to be meaningful in real-world scenarios is that the BTL model accurately represents the given pairwise comparison data. If the underlying data-generating process violates the assumptions made by the BTL model, then the model’s predictions and guarantees cannot be trusted. Hence, it is important to develop a systematic method to test the validity of the BTL assumption on real-world datasets. Such a method can help identify scenarios where a BTL model can provide a useful approximation to the underlying preferences or skills, thereby providing essential guidance for downstream applications.
In this work, our broad objective is to develop and address key questions concerning hypothesis testing for the BTL model given pairwise comparison data (also see [MakurSinghBTL]). Specifically, we aim to formulate and rigorously perform minimax analysis of a test that determines whether a given pairwise comparison dataset conforms to an underlying BTL model. Ideally, we would like to develop a test that does not require additional observations beyond the usual $k=O(1)$ comparisons per pair of items, which is known to be sufficient for consistent parameter estimation (which means the relative $\ell^{2}$-estimation error vanishes as $n\to\infty$) [Chenetal2019]. As we will see, the critical threshold for our testing problem as well as its type I and type II probabilities of error converge to zero as $n\to\infty$ with $k=O(1)$ observations. Hence, in this sense, testing for BTL models can be performed under similar conditions on $k$ and $n$ as estimation of BTL model parameters. We next delineate some related literature and then present the main contributions of this work.
@@ I-A Related Literature
This work lies at the confluence of two fields of study: preference learning and hypothesis testing. On the preference learning front, the problem of analyzing ranked preference data has become increasingly important with the rise in the availability of data on consumer preferences and web surveys. In a complementary vein, the analysis of preference data, such as pairwise comparisons, has a rich history starting from the seminal works [BradleyTerry1952, Plackett1975, Luce1959, McFadden1973, Thurstone1927, Zermelo1929]. Among the various models proposed for ranking and analyzing pairwise comparison data, the BTL model [BradleyTerry1952] has emerged as one of the most widely used and well-studied approaches. Initially introduced in [Zermelo1929] for estimating participants’ skill levels in chess tournaments, the BTL model is a special case of the more general Plackett-Luce model [Luce1959, Plackett1975], originally developed in social choice theory, mathematical psychology, and statistics contexts. The BTL model finds applications in diverse domains, such as sports [Cattelan2012, cattelan2013dynamic], psychology [Matthews2018pain], ranking of journals [Stigler1994], fine-tuning large language models [InstructGPT], etc. We refer the readers to [Diaconis1988, Shah2019] for a comprehensive overview of different models of rankings.
@@
In the literature, many studies have focused on estimating parameters of the BTL model and characterizing the corresponding error bounds, cf. [Yellott1977, SimonsYao1999, Hunter2004, NegahbanOhShah2012, NegahbanOhShah2017, Negahbanetal2018, Shahetal2016, Chenetal2019]. For example, [NegahbanOhShah2017] introduces and analyzes spectral estimation methods, [Chenetal2019] presents non-asymptotic bounds for relative $\ell^{\infty}$-estimation errors of normalized vectors of skill scores for both spectral and maximum likelihood estimators, and [Shahetal2016] provides graph-dependent $\ell^{2}$-estimation error bounds for the maximum likelihood estimator. Another emerging line of work includes uncertainty quantification for the estimated parameters, cf. [UncertaintyQuantificationBTL, LagrangianInference, UncertaintyQuantificationMultiway, UncertaintyQuantificationCovariates]. For example, [UncertaintyQuantificationBTL] proves the asymptotic normality of the maximum likelihood estimator as well as the spectral estimator for BTL skill parameters. In a different vein, [JadbabaieMakurShah2020, JadbabaieMakurShah2020journal] develop a technique for estimating the underlying skill distribution of agents participating in a tournament, introducing a Bayesian flavor to the BTL estimation problem.
@@
On the other hand, hypothesis testing also has a rich history in statistics, ranging from Pearson’s $\chi^{2}$-test [Lehmann2006] to non-parametric tests [GrettonBorgwardtRaschScholkopf2012]. The classical literature develops some tests for the IIA assumption, e.g., [HausmanMcFadden, McfaddenTyeTrain, ShevFujiiHsiehMcCowan2014], but these tests are based on asymptotic $\chi^{2}$-approximations. Moreover, testing for BTL models has been explored far less than the estimation question in recent times. And to the best of our knowledge, no study has developed rigorous hypothesis tests with minimax analysis to determine the validity of the BTL assumption in the literature. The minimax perspective of hypothesis testing, which we specialize in our setting, was initially proposed by [Ingster1994]. Among related works to ours, [RastogiBalakrishnanShahAarti2022] analyzed two-sample testing on pairwise comparison data, and [Seshadri2020] derived lower bounds for testing the IIA assumption given preference data. The formulations in [RastogiBalakrishnanShahAarti2022] and [Seshadri2020] are different to our formulation here and in [MakurSinghBTL]. For example, we quantify the separation distance of a pairwise comparison model from the class of BTL models in terms of the Frobenius norm rather than the variant of total variation (TV) distance used in [Seshadri2020]. For the special case of pairwise comparisons under a complete induced observation graph, the lower bounds in [Seshadri2020] agree with ours in terms of the high-level scaling law of the critical threshold with respect to $n$ and $k$. However, their study does not focus on providing statistical tests and corresponding upper bounds as we do in this work. Finally, in the alternative setting of online preference learning, [Haddenhorst2021] studied the testing problem for various forms of stochastic transitivity and established upper and lower bounds on the sample complexity for testing.
@@
In this work, we also briefly explore the stability of rankings made under a BTL assumption when the underlying model does not conform to this assumption. Indeed, since the BTL assumption may not perfectly hold in various scenarios, it is crucial to examine the stability of rankings obtained under the BTL assumption. In the literature, [ShahWainwright2018] provided empirical evidence that the BTL assumption is not very robust to changes in the pairwise comparison matrix, and [cucuringu2016sync] proposed the Sync-Rank algorithm, which treated the ranking problem as a group synchronization problem, thereby eliminating the dependence of the algorithm on the BTL assumption. Furthermore, recent work by [datar2022byzantine] demonstrated the ineffectiveness of standard spectral methods for BTL parameter estimation in the presence of even small fractions of Byzantine voters, and proposed a more robust Byzantine spectral ranking algorithm. In contrast to these works, we introduce and utilize a new measure of separation distance that quantifies deviation from the class of BTL models. As we will see, this measure is instrumental in our analysis to quantify the stability of BTL assumptions in the context of rankings with respect to classical Borda count rankings.
@@ I-B Main Contributions
    We define the critical threshold for our testing problem and establish an upper bound on it in Theorem 2 for general induced observation graphs (satisfying mild assumptions). Furthermore, we also derive upper bounds on the type I and type II probabilities of error in Theorem 3. These bounds provide insights into the influence of various parameters on the error decay rate.
@@
    We also prove several auxiliary results. For example, during our analysis, we obtain $\ell^{2}$-estimation error bounds for (virtual) BTL parameters in Lemma 6 when the data is actually generated by a general pairwise comparison model as opposed to a BTL model. These bounds also highlight the robustness of the spectral ranking method under model mismatch. As another example, we utilize the notion of separation distance mentioned above to analyze the stability of BTL assumptions in the context of rankings in Appendix E. More specifically, we investigate the deviation from the BTL condition that is sufficient for the ranking produced under the BTL assumption to differ from the classical Borda count ranking [borda1781]. Our results in Proposition 10 show that a deviation of $O(1/\sqrt{n})$ from the BTL condition is sufficient to produce inconsistent BTL and Borda rankings.
@@ I-C Outline
The paper is organized as follows. Several notational preliminaries are presented in Section I-D. In Section II, we introduce general pairwise comparison models, define the necessary terminology, and present the various regularity assumptions necessary for our analysis. In Section III, we present the main results of our work. Specifically, in Section III-A, we introduce a notion of separation distance that allows us to measure the deviation of a pairwise comparison model from the class of all BTL models. Additionally, we mathematically formulate the hypothesis testing problem and introduce the associated testing rule. Section III-B provides an upper bound on the scaling of the critical threshold for this hypothesis test. In Section III-C, we provide upper bounds on type I and type II probabilities of error. Moreover, Section III-D introduces a matching lower bound on the critical threshold of the test proving minimax optimality (up to constant factors). In Section III-E, we provide examples of different classes of graphs that meet the required assumptions for all pairwise comparison models. In our paper, formal proofs of various propositions are available in Section IV, while the upper bound proofs for the critical threshold (Theorem 2) and type I and type II probabilities of error (Theorem 3) are provided in Section V. The proof of our lower bound is presented in Section VI. In Section VII, we present numerical simulations for synthetic data that support our theoretical results and also apply our testing rule to both real-world and synthetic datasets. Finally, in Section VIII, we reiterate our results and provide some directions for future research. Additionally, we present our remaining proofs and characterize the stability of the BTL model in the Appendices.
I-D Notational Preliminaries
We briefly collect some notation here that is used throughout this work. We let ${\mathbb{R}}=(-\infty,\infty)$, ${\mathbb{R}}_{+}=[0,\infty)$, and ${\mathbb{N}}\triangleq\{1,2,3,\dots\}$ denote the sets of real, positive real, and natural numbers, respectively. For any $n\in{\mathbb{N}}$, we let ${\mathbb{R}}^{n}$ and ${\mathbb{R}}^{n\times n}$ be the sets of all $n$-length vectors and $n\times n$ matrices, respectively, ${\mathbf{1}}_{n}\in{\mathbb{R}}^{n}$ be the column vector with all entries equal to $1$, and $n]\triangleq\{1,\dots,n\}$. For any (index) set ${\mathcal{S}}\subseteq[n]$, we let ${\mathbf{1}}_{\mathcal{S}}\in{\mathbb{R}}^{n}$ denote a vector in which the entries are $1$ for elements in the set $\mathcal{S}$} and 0 otherwise. Moreover, we let ${\mathcal{P}}_{n}$ denote the (n1)(n-1)-dimensional probability simplex in ${\mathbb{R}}^{n}$, and ${\mathbbm{1}}_{\mathcal{A}}$ denote the indicator function on the set $\mathcal{A}$. Next, for any vector $x\in{\mathbb{R}}^{n}$, $\|x\|_{2}$ and $\|x\|_{\infty}$ denote its $\ell^{2}$- and $\ell^{\infty}$-norms, $x^{p}$ denotes the entrywise $p$th power of $x$, i.e., $x^{p}=[x_{1}^{p},x_{2}^{p},\dots,x_{n}^{p}]^{\mathrm{T}}$, and $\operatorname{{diag}}(x)\in{\mathbb{R}}^{n\times n}$ is the diagonal matrix with $x$ along its principal diagonal. For any matrix $A\in{\mathbb{R}}^{n\times n}$, $\|A\|_{2}$, $\|A\|_{\mathrm{F}}$, and $\text{Tr}(A)$ denote the standard spectral norm, Frobenius norm, and trace of $A$, respectively, $A_{ij}$ denotes the $(i,j)$th element of $A$, $A_{:,i}$ denotes the $i$th column of $A$, and $A_{j,:}$ denotes the transpose of the $j$th row of $A$ (i.e., in column vector form). Let $\sigma_{1}(A)\geq\sigma_{2}(A)\geq\dots\geq\sigma_{n}(A)$ represent the (ordered) singular values of $A\in{\mathbb{R}}^{n\times n}$. For an entrywise strictly positive vector $\pi\in{\mathbb{R}}^{n}$, we define a Hilbert space $\ell^{2}(\pi)$ on ${\mathbb{R}}^{n}$ with inner product $\langle x,y\rangle_{\pi}=\sum_{i=1}^{n}\pi_{i}x_{i}y_{i}$ and corresponding vector, operator, and weighted Frobenius norms $\|x\|_{\pi}=\sqrt{\langle x,x\rangle_{\pi}}$, $\|A\|_{\pi}=\sup_{\|x\|_{\pi}=1}\|x^{\mathrm{T}}A\|_{\pi}$, and $\|A\|_{\pi,\mathrm{F}}=\big(\sum_{i}\sum_{j}\pi_{j}A^{2}_{ij}\big)^{1/2}$, respectively. Finally, we utilize standard Bachmann-Landau asymptotic notation, e.g., $O(\cdot)$, where it is assumed that the parameter $n\rightarrow\infty$ or the parameter $k\rightarrow\infty$, where $n$ represents the number of items and $k$ is the number of pairwise comparisons per pair of items. Throughout this paper, we use “high probability” to refer to a probability of at least $1-1/\text{poly}(n)$, where $\text{poly}(n)$ captures a polynomial function of $n$.
II Formal Setup and Goal